a to the 10 power divided by a to the 7 power
how do i solve this?

Answer by ptaylor(2198)

a to the 10 power divided by a to the 7 power
how do i solve this?

a^10=a*a*a*a*a*a*a*a*a*a
---- -----------------------------
a^7= a*a*a*a*a*a*a

Now we can start cancelling a's and when we are finished, we will have a*a*a in the numerator so our answer is:
a^10
----=a^(10-7)=a^3
a^7
When you divide, you subtract exponents. When you multiply, you add them.

Hope this helps----ptaylor